# The Basis of Foot-Sortable Sock Orderings

### v2

@@ 2.3. Algorithm
    Otherwise, we recursively run the main algorithm on the smaller ordering $X-d$.
@@
    If it fails, then by our recursive assumption $X-d$ is unsortable, so $X$ is unsortable.
@@
    Otherwise, $X-d$ is sortable and we have the following two possibilities.
@@
      When we attempt to sort $X$ using the same sequence of colors that fully sorted $X-d$, there is a stage in which the color $d$ only appears as the top color of the stack, which means that we can fully sort $X$ by sorting $d$ at this stage and then continuing with the sequence that was used to sort $X-d$.
@@
We now present the main algorithm. It closely follows the proof Theorem 3 in Section 2.5, so the proof of Theorem 3 also serves as a proof that this algorithm is correct.

### v3

@@ Appendix A Algorithm
    Otherwise, we recursively run the main algorithm on the smaller ordering XdX-d.
@@
    If it fails, then by our recursive assumption XdX-d is unsortable, so $X$ is unsortable.
@@
    Otherwise, XdX-d is sortable and we have the following two possibilities.
@@
      When we attempt to sort $X$ using the same sequence of colors that fully sorted XdX-d, there is a stage in which the color $d$ only appears as the top color of the stack, which means that we can fully sort $X$ by sorting $d$ at this stage and then continuing with the sequence that was used to sort XdX-d.
@@
We now present the main algorithm. It closely follows the proof Theorem 2.